OpenStudy (anonymous):

solve for log x +log(x-21)=2

OpenStudy (raden):

simplify for left side by using the property of log, log x +log(x-21) = log x(x-21)

OpenStudy (raden):

setting 2 be log 100

OpenStudy (raden):

now u have an equation : log x(x-21) = log 100 cancel the log from both sides, x(x-21) = 100 x^2 - 21x - 100 = 0 solve for x

OpenStudy (raden):

@hperkins u can use by factoring or complete squares, also using the quadratic formula